OpenStudy (anonymous):

i have the NxN matrix A=[ a 1 0 . . . . . 0 1 a 1 . . . .... 0 . . . . . . . .. 0 0 0 . . 1 a 1 0 0 0 . . 0 1 a ] and i want the |A|. I have |An|=a|An-1| - |An-2|. can i find some more specific ?

OpenStudy (cristiann):

This is a second-order linear recurrence (a linear difference equation) First, you have to calculate several beginning terms (A1, A2, A3, A4, A5, etc... should be enough ...:) ) Then, you have to suppose for An the form: An=c*x^n+d*y^n [sum of two geometric progressions with coefficients] When you replace you get that x and y are solutions of the characteristic equation t^2-at+1=0; Depending on a, you have a discussion with three cases: solutions real, distinct (you find x and y, then c and d by using x and y and the terms A1 and A2 - the results are a bit ugly) solutions real, equal (you use the modified general form An=c*x^n+d*n*x^n) and you should get: for a=2: An=n+1 for a=-2: An=(n+1)(-1)^n solutions complex: since x and y are complex roots of a real second-order equation, they complex conjugates (and they have the complex modulus 1 and the same argument arg - the (polar) angle from their trigonometric form). So you should again change the form you are looking for into: An=c*sin(arg)+d*cos(arg) You may find the general theoretic details for the above in a book about (linear) difference equations (or recurrent relations). If you need some info, please let me know... :)